### v1

# Reinforcement Learning with Intrinsically Motivated Feedback Graph for Lost-sales Inventory Control

[…]
Reinforcement Learning with Intrinsically Motivated Feedback Graph for Lost-sales Inventory Control
[…]
eejzhang@ust.hk
[…]
Reinforcement learning (RL) has proven to be well-performed and general-purpose in the inventory control (IC). However, further improvement of RL algorithms in the IC domain is impeded due to two limitations of online experience. First, online experience is expensive to acquire in real-world applications. With the low sample efficiency nature of RL algorithms, it would take extensive time to train the RL policy to convergence. Second, online experience may not reflect the true demand due to the lost sales phenomenon typical in IC, which makes the learning process more challenging. To address the above challenges, we propose a decision framework that combines reinforcement learning with feedback graph (RLFG) and intrinsically motivated exploration (IME) to boost sample efficiency. In particular, we first take advantage of the inherent properties of lost-sales IC problems and design the feedback graph (FG) specially for lost-sales IC problems to generate abundant side experiences aid RL updates. Then we conduct a rigorous theoretical analysis of how the designed FG reduces the sample complexity of RL methods. Based on the theoretical insights, we design an intrinsic reward to direct the RL agent to explore to the state-action space with more side experiences, further exploiting FG’s power. Experimental results demonstrate that our method greatly improves the sample efficiency of applying RL in IC. Our code is available at https://anonymous.4open.science/r/RLIMFG4IC-811D/
1 Introduction
Inventory control (IC) is a crucial and practical problem, serving as the basis of business efficiency for supply chain management. IC problems are challenging due to the intractability caused by the inherent complexity and the difficulty of finding optimal solutions within a reasonable timeframe. In the past few decades, researchers Zipkin (2008a), Bai et al. (2023), Xin (2021), and Chen et al. (2024) have designed a few heuristic methods based on specific model assumptions. However, such methods often face challenges due to the curse of dimensionality (Goldberg et al. 2021), where the problem size grows exponentially as the lead time increases. Here lead time refers to the duration between placing and receiving an order. Another limitation is that these model-based methods are not flexible enough to generalize to various environmental settings. These limitations call for a more adaptable approach to IC problems, for which data-driven control methods stand out as a promising alternative.
Reinforcement learning (RL) has gained significant attention as a powerful data-driven technique for solving complex sequential decision-making problems. In particular, it offers several advantages for addressing the challenges of IC problems. Firstly, RL allows for the discovery of optimal policies without relying on strong problem-specific assumptions, enabling more generalizable solutions (Nian et al. 2020). Secondly, when combined with the deep neural network, Deep RL (DRL) can handle large state and action spaces, making it suitable for problems with high-dimensional state variables (Dehaybe et al. 2024). Early works demonstrate the feasibility of RL in IC problems. Oroojlooyjadid et al. (2022) showcases the ability of a Deep Q-network (DQN) to discover near-optimal solutions for the widely recognized beer distribution game. Gijsbrechts et al. (2022) introduces the A3C algorithm into IC problems and show that it can achieve acceptable performance but not better than heuristic methods. Stranieri and Stella (2023) benchmarks various DRL methods such as A3C, PPO, and vanilla policy gradient (VPG) in IC problems. More recently, researchers tend to explore how DRL can address various scenarios within the IC problems. These scenarios include: non-stationary uncertain demand (Dehaybe et al. 2024, Park et al. 2023), multi-product (Sultana et al. 2020, Selukar et al. 2022), variable kinds of products (Meisheri et al. 2020, 2022), multi-echelon supply chains (Wu et al. 2023, Stranieri et al. 2024), one-warehouse multi-retailer (Kaynov et al. 2024), and the stochastic capacitated lot sizing problem (Van Hezewijk et al. 2023).
However, traditional RL methods are characterized by low sample efficiency, which becomes a significant barrier when implementing these techniques in real-world IC scenarios, as obtaining experiences can be both costly and time-consuming (Boute et al. 2022, De Moor et al. 2022). Furthermore, this sample inefficiency issue is enlarged in lost-sales IC problems because of censored demands (Chen et al. 2024), which refers to the phenomenon when customers’ real demands are unobservable due to insufficient inventory. For instance, if the order is placed daily, it takes over a year to generate four hundred experiences and part of them may be censored, making them too few to update the RL policy. De Moor et al. (2022) tries to alleviate this problem by incorporating heuristic knowledge, such as the base-stock policy, into DQN with reward-shaping. Although this method can improve the sample efficiency, it still relies on specific model heuristics, making it hard to generalize to different scenarios. Overall, resolving the low sample efficiency of RL without strong heuristics is crucial in solving IC (especially lost-sales IC) problems. The detailed related work is in Appendix 6.
By addressing the above-mentioned limitation, this paper proposes a novel decision framework that combines the reinforcement learning with feedback graphs (RLFG) and intrinsically motivated exploration (IME):
1) We tailor the feedback graph (FG) based the general property of lost-sales IC problems rather than strong heuristics (e.g. known demand distribution). In particular, the connectivity of FG is adjusted dynamically based on the relationship between the demand and inventory rather than being static in the environment. With FG, the sample efficiency in the training process is significantly improved by allowing the agent to acquire not only online experiences but also side experiences from FG.
2) We conduct a theoretical analysis of how FG reduces the sample complexity with Q-learning as an example. It demonstrates that FG decreases the sample complexity by improving the update probabilities across all state-action pairs.
3) Inspired by these theoretical insights, we design a novel intrinsic reward that guides the RL algorithm to explore towards the state-action space where more side experiences can be obtained thereby further boosting sample efficiency.
4) We evaluate our method on the standard discrete lost-sales inventory control environment. Our empirical results demonstrate the superior sample efficiency improvement by FG and the intrinsic reward separately, underlining the effectiveness of our design.
2 Background and Problem Formulation
2.1 Reinforcement Learning with Feedback Graph
RLFG is proposed by Dann et al. (2020) to reduce the sample complexity of RL. In typical RL scenarios, an agent can only get one experience each time as feedback that can be used to update the policy. However, when some prior knowledge about the environments is available, it becomes possible to observe additional experiences involving other states and actions. RLFG aims to explore how RL algorithms can benefit from these side experiences by constructing a feedback graph (FG). Here “side experiences” bears the same meaning as “side observations” in Dann et al. (2020), and we use the term “side experiences” to avoid confusion with “observations” in partially observed MDPs. FG is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the vertex set $\mathcal{V}=\{\bm{v}|\bm{v}=(\bm{s},a)\}$ and $\mathcal{E}$ is the edge set =$\mathcal{E}=\{\bm{v}\rightarrow\bm{\bar{v}}\}$, formalized by the side information indicating that if the agent visits $\bm{v}$, it can also observe other vertices $\bm{\bar{v}}$. The total experiences $\mathcal{O}_{t}$ observed by the agent at $t$ from $\mathcal{G}$ is
$\mathcal{O}_{t}(\mathcal{G})=\{(\bm{s}_{t},a_{t},r_{t},\bm{s}_{t+1})\}\cup\{\bm{\bar{s}_{t}},\bar{a_{t}},\bar{r_{t}},\bm{\bar{s}}_{t+1}\}.$ (1)
2.2 Lost Sales Inventory Control Problem
Table 1: The demands and state variables.
Symbol Representation Symbol Representation
$L\geq 0$ Order lead time $a_{t}\in\mathbb{N}$ Order at time $t$
$t=1,...,T$ Time index $a^{\text{max}}\in\mathbb{N}$ Maximum order
$d_{t}\in\mathbb{N}$ Real demand in period $t$ $c\in\mathbb{R}$ Unit cost of procurement
$d^{o}_{t}\in\mathbb{N}$ Demand observed in period $t$ $h\in\mathbb{R}$ Unit cost of holding inventory
$d^{\text{max}}\in\mathbb{N}$ Maximum demand $p\in\mathbb{R}$ Unit cost penalty of lost sales
$\bm{s}_{t}\in\mathbb{N}^{L}$ State: $(y_{t},a_{t+1-L},...,a_{t-1})$ $\gamma\in(0,1]$ Discount factor
$y_{t}\in\mathbb{N}$ Inventory at time $t$, $y^{\text{max}}\in\mathbb{N}$ Maximum inventory at time $t$,
after the order due at $t$ arrives after the order due at $t$ arrives
We formulate a standard, single-item, discrete-time, and lost-sales IC problem following Zipkin (2008a), Gijsbrechts et al. (2022) and Xin (2021) with environment variables in Table 1. The “lost-sales” means that customers leave if there is not enough inventory without any way to record the excess demand. The lost-sales IC problem considers three kinds of costs in each time step $t$, which are the cost of procurement $f_{1}(\bm{s}_{t})$, the cost of holding inventory $f_{2}(\bm{s}_{t})$, and the cost penalty of lost sales $f_{3}(\bm{s}_{t})$. The cost terms are specified below, where $[x]^{+}=\max(x,0)$.
$f_{1}(\bm{s}_{t},a_{t})=ca_{t},f_{2}(\bm{s}_{t})=h[y_{t}-d_{t}]^{+},f_{3}(\bm{s}_{t})=p[d_{t}-y_{t}]^{+}.$ (2)
In this problem, the real demand is unobservable when it exceeds the current inventory. Thus we define the real demand $d_{t}$ as a random variable and the observed demand $d_{t}^{o}$, which may be censored by $y_{t}$ and $d_{t}$, given as
$d_{t}^{o}=\min(d_{t},y_{t})=\begin{cases}d_{t},&\ \text{if}\ d_{t}\leq y_{t}\\ y_{t},&\ \text{otherwise}\end{cases}.$ (3)
Here, the term “censored” indicates the otherwise case in Equation 3, when we can only observe all inventory is sold but do not know the value of $d_{t}$. The objective is to minimize the total cost over the time horizon $T$ under the uncertainty on the demand side, which is
$\min_{\{a_{t}|t=0,...,T\}}\sum_{t=0}^{T}\gamma^{T-t}f(\bm{s}_{t},a_{t})=\sum_{t=0}^{T}\gamma^{T-t}[f_{1}(\bm{s}_{t},a_{t})+f_{2}(\bm{s}_{t})+f_{3}(\bm{s}_{t})].$ (4)
2.3 MDP Formulation
To better understand the lost-sales IC problem from the perspective of RL, we formulate it into an infinite-horizon MDP with discounted rewards (Li et al. 2020), represented by $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$. The MDP consists of the state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(\bm{s}^{\prime}|\bm{s},a)$, reward function $R$, and discount factor $\gamma$. The detailed composition of MDP is shown as follows:
State: The state includes the inventory at time $t$ after receiving the orders and all upcoming orders due to lead time. We define it as $\bm{s}_{t}=(y_{t},a_{t+1-L},...,a_{t-1})$.
Action: The action is the amount to be ordered for future sales, given as $a_{t}=\{0\leq a_{t}\leq a^{\text{max}},a_{t}\in\mathbb{N}\}$, where $a^{\text{max}}$ is the maximum amount that can be ordered for this item.
Reward: Since the goal is to minimize the cumulative discounted cost, the reward at each time step $t$ is defined as the opposite number of costs, which is $r_{t}=R(\bm{s}_{t},a_{t})=-f(\bm{s}_{t},a_{t})$.
Transition Function: The transition function is defined as $\bm{s}_{t}=(y_{t},a_{t+1-L},...,a_{t-1})\to\bm{s}_{t+1}=(y_{t+1},a_{t+2-L},...,a_{t})$. The inventory transition is defined as $y_{t+1}=[y_{t}-d_{t}]^{+}+a_{t+1-L}$ due to the lead time. As $a_{t+1-L}$ has been received, $a_{t+1+i-L}$ in $\bm{s}_{t+1}$ will replace $a_{t+i-L}$ in $\bm{s}_{t}$ for $i=1,...,L-1$.
3 Method
Figure 1: IC decision framework based on RLFG and IME. Solid lines depict sampling steps and dashed lines represent model-updating steps.
3.1 IC Decision Framework with RLFG and IME
The proposed decision framework for IC problems aims to enhance the sample efficiency of off-policy RL algorithms by integrating RLFG and IME together. This framework relies on two key assumptions. First, the real demand is unknown and the observed demand may be censored. Second, the experiences generated in real-world operations are limited due to the cost involved in collecting online data. As illustrated in Figure 1, the FG module generates side experiences to improve the sample efficiency and the intrinsic reward module aids the exploration to further exploit the power of the FG module. Here, the RL module can be any off-policy RL algorithm, such as DQN, DDPG, Rainbow, or TD3. Note that experiences from both reply buffers are sampled with no differences in the model-updating step. Then a new set of side experiences are generated based on the sampled experiences. These new side experiences are just used to calculate intrinsic rewards of the sampled experiences. Here we provide Rainbow-FG as an example in Algorithm 1.
Algorithm 1 Rainbow-FG Algorithm.
1:  Initialize the distributional Q-network $Q$ with parameter $\theta$, the target Q-network $Q^{\prime}$ with parameter $\phi$, the replay buffer $\mathcal{R}$, and the side replay buffer $\mathcal{R}^{s}$
2:  for episode = $1,...,M$ do
3:     Initialize the environment state $s_{0}$
4:     for $t=0,...,T$ do
5:        Get the order $a_{t}=\begin{cases}\text{argmax}_{a}\ Q(\bm{s}_{t},a),&\ w.p\ \ 1-\epsilon\\ \text{a random action},&\ w.p\ \ \epsilon\end{cases}$
6:        Execute action $a_{t}$ and get next state $\bm{s}_{t+1}$ and reward $r_{t}$
7:        Get the observed demand $d_{t}^{o}$
8:        Get the side experiences $\{(\hat{\bm{s}_{t}},\hat{a_{t}},\hat{r_{t}},\hat{s}_{t+1},d^{o}_{t})\}\to$ Algorithm 2
9:        $\mathcal{R}=\mathcal{R}\cup(\bm{s}_{t},a_{t},r_{t},\bm{s}_{t+1},d_{t}^{o})$ and $\mathcal{R}^{s}=\mathcal{R}^{s}\cup\{(\hat{\bm{s}_{t}},\hat{a_{t}},\hat{r_{t}},\hat{\bm{s}}_{t+1},d_{t}^{o})\}$
10:        Model Update:
11:        Randomly sample a batch of experiences $\{(\bm{s},a,r,\bm{s}^{\prime},d^{o})\}$ from $\mathcal{R}\cup\mathcal{R}^{s}$
12:        Get all side experiences $\{(\hat{\bm{s}},\hat{a},\hat{r},\hat{\bm{s}}^{\prime},d^{o})\}$ for each experience in $\{(\bm{s},a,r,\bm{s}^{\prime},d^{o})\}\to$ Algorithm 2
13:        Get the intrinsic reward $r^{in}$ of this batch of experiences $\to$ Algorithm 3
14:        Get the final reward $r=(1-\beta)\times r+\beta\times r^{in}$
15:        Update parameters
3.2 Feedback Graph in Inventory Control
Motivated by the reduction in sample complexity by FG (Dann et al. 2020), we incorporate FG into the IC problem. FG is naturally suitable for the IC problem since the IC problem is a structured MDP with most environmental transition components determined and predictable depending on the demand. Furthermore, the demand is usually independent of the state-action space. If we can observe the demand, the experiences of other state-action pairs can also be obtained through leveraging the underlying properties of such structured MDPs. Thus the main challenge is how to construct FG considering the lost-sales property, whose demand is potentially censored. To solve this problem, we propose to construct FG dynamically based on the observed demand. If the observed demand is real, FG generates side experiences of all other state-action pairs. Even if the observed demand is censored, FG can generate side experiences of the state-action pairs whose inventory is less than the observed demand.
Algorithm 2 Feedback Graph Module.
1:  Input: experience $(\bm{s}_{t}=(y_{t},a_{t+1-L},...,a_{t-1}),a_{t},r_{t},\bm{s}_{t+1},d_{t}^{o})$
2:  Initialize a temporary buffer ${}\mathcal{B}=\{\}$
3:  if $d_{t}^{o}=y_{t}$ (censored case) then
4:     $y^{\text{bound}}=y_{t}$
5:  else
6:     $y^{\text{bound}}=y^{\text{max}}$
7:  for $\hat{y}_{t}=0,...,y^{\text{bound}}$ do
8:     for each term of $(\hat{a}_{t+1-L},...,\hat{a}_{t-1})=0,...,a^{\text{max}}$ do
9:        $\hat{s}_{t}=(\hat{y}_{t},\hat{a}_{t+1-L},...,\hat{a}_{t-1})$
10:        for $\hat{a}_{t}=0,...,a^{\text{max}}$ do
11:           Get $\hat{r}_{t}$ and $\hat{\bm{s}}_{t+1}$
12:           $\mathcal{B}=\mathcal{B}\cup(\hat{\bm{s}_{t}},\hat{a_{t}},\hat{r_{t}},\hat{\bm{s}}_{t+1},d_{t}^{o})$
13:  Return $\mathcal{B}$
Algorithm 2 shows the details of the FG module. When $d^{o}_{t}$ is uncensored, FG is a complete graph and all state-action pairs can be used to generate side experiences based on $d^{o}_{t}$. When $d^{o}_{t}$ is censored, it becomes more complex. For the state-action pairs having larger inventory numbers than that in $\bm{s}_{t}$, $d^{o}_{t}$ is not a correct demand for them and unfortunately $d_{t}$ is unknown. Still using $d^{o}_{t}$ will generate wrong side experiences. We can only use the state-action pairs having smaller inventory numbers than that in $\bm{s}_{t}$ to generate side experiences, which means that FG is a partially connected graph. Note that FG is still dynamically constructed based on $d^{o}_{t}$ under the censored case since different $d^{o}_{t}$ generates different numbers of side experiences.
3.3 Theoretical Analysis
We conduct a quantitative analysis of how the sample complexity is reduced with the FG in the lost-sales IC environment. The qualitative analysis based on the property of RLFG is in Appendix 7. Here we provide an analysis based on Q-learning for simplicity. The detailed proof is in Appendix 8
Without loss of generality, we restrict some definitions in section 2.3 and define some new concepts. We consider the reward function $R:\mathcal{S}\times\mathcal{A}\to(0,1)$. The demand $d_{t}\sim P_{d}(d)$ can obey any independent discrete distribution and $d^{\text{max}}<y^{\text{max}}$. We define $\pi_{b}$ as the stationary behavior policy and $\mu(\bm{s},a)$ as the stationary distribution of the Markov chain under $\pi_{b}$ and $P(\bm{s}^{\prime}|\bm{s},a)$, which is the same as the update probability in typical RL. In RLFG, the stationary distribution does not change but the update probability becomes $\tilde{\mu}(\bm{s},a)\geq\mu(\bm{s},a)$ because of the side experiences.
Scenario 1: Consider a graph $\mathcal{G}$ with all state-action pairs as the nodes and there is no edge between the nodes. It can be regarded as $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$, where $\mathcal{G}_{1}=\emptyset$ and $\mathcal{G}_{2}$ consists of nodes without any edge.
Lemma 1: The sample complexity of the asynchronous Q-learning under scenario 1 is analyzed in Li et al. (2020), which is $\tilde{O}(\frac{1}{\mu_{\text{min}}(1-\gamma)^{5}\epsilon^{2}}+\frac{t_{\text{mix}}}{\mu_{\text{min}}(1-\gamma)})$, where $\mu_{\text{min}}=\min_{(s,a)\in\mathcal{G}}\mu(\bm{s},a)$ and $t_{\text{mix}}$ is the mixing time of the chain.
Lemma 1 indicates that the sample complexity of Q-learning without FG is determined by $\mu_{\text{min}}$. We will show how Q-learning with FG improves $\mu(\bm{s},a)$ to reduce the sample complexity.
Scenario 2: Consider a graph $\mathcal{G}$ with all state-action pairs as the nodes. Assume for nodes satisfying $y\geq d_{t}$, once one node is sampled, all of these nodes can be sampled and updated simultaneously. For nodes with $y<d_{t}$, only nodes with $y^{\prime}\leq y$ can be sampled and updated simultaneously. Thus $\mathcal{G}$ can be regarded as $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$.
Theorem 1: The update probability for Q-learning with FG under scenario 2 is:
$\tilde{\mu}(\bm{s},a)=\underbrace{\sum_{d=0}^{d^{\text{max}}}P_{d}(d)\sum_{\begin{subarray}{c}(\bm{\bar{s}},\bar{a})\in\mathcal{G}\\ \bar{y}\geq d\end{subarray}}\mu(\bm{\bar{s}},\bar{a}|d)}_{\text{Uncensored term}}+\underbrace{\sum_{d=0}^{y-1}P_{d}(d)\sum_{\begin{subarray}{c}(\bm{\bar{s}},\bar{a})\in\mathcal{G}\\ y\leq\bar{y}\leq d\end{subarray}}\mu(\bm{\bar{s}},\bar{a}|d)}_{\text{Censored term}}.$ (5)
Conclusion: We can obtain the relationship between $\tilde{\mu}(\bm{s},a)$ and $\mu(\bm{s},a)$ in Equation (6). It shows that FG not only improves $\mu_{\text{min}}$ but also improves that of all the state-action pairs.
$\tilde{\mu}(\bm{s},a)=\mathbb{E}_{d\sim P_{d}}[\tilde{\mu}(\bm{s},a|d)]\geq\mathbb{E}_{d\sim P_{d}}[\mu(\bm{s},a|d)]=\mu(\bm{s},a).$ (6)
3.4 Intrinsically Motivated Exploration
To further utilize the benefits of FG, we design an intrinsic reward by incorporating the information of the side experiences into the curiosity-driven exploration. In particular, for a state-action pair, the number and the average uncertainty of the side experiences generated by this state-action pair are incorporated into the intrinsic reward. With this intrinsic reward, the agent is directed towards the state-action space where more side experiences can be generated.
This idea is inspired by the theoretical analysis based on Equation 5. The uncensored term indicates how much the uncensored case contributes to improving the probability of being updated. This term contributes to all state-action pairs so improving this term can improve the probability of being updated for all state-action pairs. The censored term indicates how much benefit the current state-action pair can obtain from the censored cases. For the censored case, the sample complexity can be further reduced by visiting the state-action pairs with larger inventory. To satisfy both conditions, designing a behavior policy manually is difficult and not general enough. Thus we design an intrinsic reward, written in Equation 7 because both conditions lead to generating more side observations.
$r^{in}_{i}=r^{in}_{i}+log_{10}(J)\times\frac{1}{J}\sum_{j=1}^{J}r^{in}_{i,j}.$ (7)
Algorithm 3 shows the details of the intrinsic reward module. We utilize the M-head DQN (Nikolov et al. 2018) with each head trained by different mini-batch of experiences to get the prediction error as curiosity rewards. The final intrinsic reward, denoted in Equation 7, consists of the curiosity reward of the experience itself ($r^{in}_{i}$) and the averaged curiosity reward of all corresponding side experiences ($r^{in}_{i,j}$) generated by this experience. The averaged value preserves the advantage of scale invariance but loses the quantity information compared with the sum value. To balance these two aspects, we multiply the averaged value with $\log_{10}(J)$ to incorporate more quantity information into the intrinsic reward. In the censored condition, where $J=y_{t}$, a larger $J$ increases the intrinsic reward, which means more side experiences. In the uncensored condition, $J=y^{\text{max}}$ is larger than $y_{t}$, which means that the intrinsic reward is more likely to be larger than that in the censored condition.
Algorithm 3 Intrinsic Reward Module.
1:  Input: Mini-batch experiences $\{(\bm{s},a,r,\bm{s}^{\prime},d^{o})\}$ and all side experiences $\{(\hat{\bm{s}},\hat{a},\hat{r},\hat{\bm{s}}^{\prime},d^{o})\}$
2:  for each experience $(\bm{s}_{i},a_{i},r_{i},\bm{s}^{\prime}_{i},d^{o}_{i})$ in $\{(\bm{s},a,r,\bm{s}^{\prime},d^{o})\}$ do
3:     $\bar{Q}^{dqn}=\frac{1}{M}\sum_{m=1}^{M}Q^{dqn}_{m}(\bm{s}_{i},a_{i})$
4:     $r^{in}_{i}=\frac{1}{M}\sqrt{\sum_{m=1}^{M}[Q^{dqn}_{m}(\bm{s}_{i},a_{i})-\bar{Q}^{dqn}]^{2}}$
5:     $\{(\hat{\bm{s}}_{i},\hat{a}_{i},\hat{r}_{i},\hat{\bm{s}}^{\prime}_{i},d^{o}_{i})\}$ are all side experiences generated from $(\bm{s}_{i},a_{i},r_{i},\bm{s}^{\prime}_{i},d^{o}_{i})$
6:     $J$ is the size of $\{(\hat{\bm{s}}_{i},\hat{a}_{i},\hat{r}_{i},\hat{\bm{s}}^{\prime}_{i},d^{o}_{i})\}$
7:     for each side experience $(\hat{\bm{s}}_{i,j},\hat{a}_{i,j},\hat{r}_{i,j},\hat{\bm{s}}^{\prime}_{i,j},d^{o}_{i})$ in $\{(\hat{\bm{s}}_{i},\hat{a}_{i},\hat{r}_{i},\hat{\bm{s}}^{\prime}_{i},d^{o}_{i})\}$ do
8:        $\bar{Q}^{dqn}=\frac{1}{M}\sum_{m=1}^{M}Q^{dqn}_{m}(\hat{\bm{s}}_{i,j},\hat{a}_{i,j})$
9:        $r^{in}_{i,j}=\frac{1}{M}\sqrt{\sum_{m=1}^{M}[Q^{dqn}_{m}(\hat{\bm{s}}_{i,j},\hat{a}_{i,j})-\bar{Q}^{dqn}]^{2}}$
10:     $r^{in}_{i}=r^{in}_{i}+log_{10}(J)\times\frac{1}{J}\sum_{j=1}^{J}r^{in}_{i,j}$
11:     Store $r^{in}_{i}$ into $(\bm{s}_{i},a_{i},r_{i},\bm{s}^{\prime}_{i},d^{o}_{i})$
12:  Return $\{(\bm{s},a,r,\bm{s}^{\prime},d^{o},r^{in})\}$
4 Experiment
To verify the performance and sample efficiency of our decision framework, we utilize the example algorithm in Algorithm 1 and test it based on the standard lost-sales IC environment. All experiments are averaged with 20 random seeds with shaded areas representing the standard deviation.
4.1 Setup
4.1.1 Baseline
Heuristic Methods: We compare our method with widely recognized and well-performed heuristic methods. All parameters used in these methods are searched in detail and the best results are selected.
  • Constant Order: The order is always a constant ($a_{t}=r^{h}$), where $r^{h}$ is a parameter.
  • Myopic 1-period (Morton 1971): Assume the distribution of $d_{t}$ is known. This method aims to minimize the expected cost at $t+L$. The order is $a_{t}=\min_{P(u_{t+L}<0)\leq\frac{c+h}{p+h}}a_{t}$, $0\leq a_{t}\leq a^{\text{max}}$.
  • Myopic 2-period (Zipkin 2008a): Since Myopic 1-period only considers the expected cost for one time step, it can be improved by considering 2-time steps.
  • Base-Stock method (Zipkin 2008a): The Base-Stock method aims to keep the inventory level constant, including upcoming orders. The order is $a_{t}=(S^{h}-\bm{1}\cdot\bm{s}_{t})^{+}$, where $S^{h}$ is a parameter.
  • Capped Base-Stock method (Xin 2021): It combines Base-Stock and Constant Order method. The order is $a_{t}=\min[(S^{h}-\bm{1}\cdot\bm{s}_{t})^{+},r^{h}]$, where $S^{h}$ and $r^{h}$ are parameters.
  • Bracket method (Bai et al. 2023): A variant of the Constant Order method. The order is $a_{t}=\lfloor(t+1)r^{h}+\theta^{h}\rfloor-\lceil tr^{h}+\theta^{h}\rceil$, where $r^{h}$ and $\theta^{h}$ are parameters.
Deep Reinforcement Learning Methods: We choose to compare Rainbow-FG with the A3C method in Gijsbrechts et al. (2022), which is an on-policy RL method; Rainbow, which is the basis of Rainbow-FG; Rainbow-FG(H), which utilizes the heuristic knowledge to continue finetuning Rainbow-FG. Rainbow-FG(H) can be regarded as the upper bound of Rainbow-FG. Since A3C is on-policy, it is not applicable to compare sample efficiency with off-policy Rainbow-FG. Thus A3C is only used to compare the optimal results.
Table 2: Parameters of the environment. $L$ and $p$ are two important parameters where $L$ reflects the dimensionality of the state and $p$ reflects the seriousness of lost sales.
Parameter $L$ $p$ $y^{\text{max}}$ $a^{\text{max}}$ $c$ $h$ $d^{m}$ $d^{\text{max}}$
Value $4$ $4$ $100$ $20$ 00 $1$ $5$ $20$
4.1.2 Environments
We test our method and baselines across different settings of the standard discrete lost-sales IC environment with demand obeying Poisson distribution, where $d^{m}$ indicates the mean value. The parameters of the environment setting are shown in Table 2 and the hyperparameters of our method are in Appendix 9. As for parameters in Table 2, we follow the common recognized settings according to Zipkin (2008a), Xin (2021), and Gijsbrechts et al. (2022). In the following sections, default parameters are used unless stated otherwise. The hyperparameter analysis is demonstrated in Appendix 12.
Table 3: Optimal result comparison. For each method, the first row is the optimal average cost and the second row is the optimality gap. The bold values indicates the values of the best four methods.
Method p=4 p=9
L=2 L=3 L=4 L=2 L=3 L=4
Optimal 4.40 4.60 4.73 6.09 6.53 6.84
Constant Order 5.27 5.27 5.27 10.27 10.27 10.27
19.8% 14.6% 11.4% 68.6% 57.3% 50.1%
Bracket 5.00 5.01 5.02 8.02 8.02 8.03
13.6% 8.9% 6.1% 31.7% 22.8% 17.3%
Myopic 1-period 4.56 4.84 5.06 6.22 6.80 7.20
3.7% 5.3% 7.1% 2.1% 4.1% 5.3%
Myopic 2-period 4.41 4.64 4.82 6.10 6.57 6.92
0.2% 0.8% 1.9% 0.2% 0.6% 1.2%
Base-Stock 4.64 4.98 5.20 6.32 6.86 7.27
5.5% 8.2% 9.9% 3.7% 5.1% 6.4%
Capped Base-Stock 4.41 4.63 4.80 6.12 6.62 6.91
0.2% 0.7% 1.5% 0.5% 1.4% 1.0%
A3C 4.54 4.74 5.05 6.38 6.73 7.07
3.2% 3.0% 6.7% 4.8% 3.1% 3.4%
Rainbow 4.52 4.72 4.91 6.28 6.72 7.09
2.7% 2.6% 3.8% 3.1% 2.9% 3.7%
Rainbow-FG (our) 4.48 4.67 4.87 6.22 6.73 6.99
1.8% 1.5% 2.9% 2.1% 3.1% 2.2%
Rainbow-FG (H) (our) 4.404 4.61 4.775 6.12 6.60 6.92
0.1% 0.2% 0.9% 0.5% 1.1% 1.2%
4.2 Optimal Results Comparison
Table 3 presents the average results and optimality gap compared to the optimal results. Without strong heuristics, Rainbow demonstrates similar performance to the A3C method in Gijsbrechts et al. (2022), whereas Rainbow-FG achieves superior results. This notable improvement can be attributed to Rainbow-FG’s ability to leverage a broader range of side experiences from the feedback graph, enabling it to converge towards better solutions compared with Rainbow.
When contrasting Rainbow-FG with the heuristic methods, Rainbow-FG still falls short of surpassing the top-performing heuristic methods, such as Myopic 2-period and Capped Base-Stock. However, Myopic 2-period method assumes the distribution of the demand is known, which is a strong assumption, and the Capped Base-Stock method needs extensive parameter searches to attain this optimal result. These parameters vary across different scenarios without patterns. Conversely, Rainbow-FG consistently achieves close performance and possesses adaptability to diverse settings without these strong heuristics. Furthermore, after incorporating the heuristic knowledge, Rainbow-FG(H) attains equivalent or even superior performance compared to the best heuristic methods, particularly in certain settings. This result indicates that Rainbow-FG has the potential to perform better than heuristic methods.
((a)) $\epsilon=0$
((b)) $\epsilon=0.1$
((c)) $\epsilon=0.4$
Figure 2: Learning process of Rainbow and Rainbow-FG under the IC environment. “Final” denotes the optimal result of Rainbow-FG. The first row is the initial-stage view and the second row is the near-convergence view of the learning process.
4.3 Sample Efficiency Comparison for FG
Figure 2 illustrates the learning process of Rainbow and Rainbow-FG with different exploration parameters. The utilization of FG significantly enhances the sample efficiency. During the initial stage, Rainbow-FG with all $\epsilon$ consistently outperforms Rainbow. Towards the end, Rainbow-FG with $\epsilon=0$ converges to the final result first and subsequently trails Rainbow-FG with $\epsilon=0.1$. Conversely, Rainbow fails to converge to the final result throughout the entire 100 episodes. Furthermore, FG contributes to improving the stability of the learning process. Notably, the learning process of Rainbow-FG exhibits lower standard deviation and more stable learning curves. Under low exploration conditions ($\epsilon=0$), Rainbow-FG demonstrates significantly faster learning compared to Rainbow. This outcome signifies the benefits of FG when exploration is not favorable. Experiments on more settings are shown in Appendix 11.
4.4 Sample Efficiency Comparison for Intrinsic Reward
This section investigates the effect of intrinsic reward designed for the IC problem. Figure 3 illustrates the learning process of Rainbow-FG with and without the intrinsic reward, referred to as Rainbow-FG w/ inr and Rainbow-FG w/o inr, respectively. The results demonstrate that the designed intrinsic reward significantly improves sample efficiency during the initial stages of training in the environment with $p=4,L=4$, and $d^{m}=5$. However, as the learning process reaches the Constant Order level, the intrinsic reward does not exhibit a substantial impact. We attribute it to the simplicity of the environment, which weakens the effect of the intrinsic reward. To further evaluate its effectiveness, we conduct experiments in more complex settings with $p=19,L=8$, and varying values of $d^{m}=$ (5, 10, 15). In these environments, we observe a clear effect of the intrinsic reward during the whole training process. Moreover, Rainbow-FG w/ inr demonstrates a more stable training process compared to Rainbow-FG w/o inr, achieving slightly lower cost.
((a)) p=4, L=4, $d^{m}$=5
((b)) p=19, L=8, $d^{m}$=5
((c)) p=19, L=8, $d^{m}$=10
((d)) p=19, L=8, $d^{m}$=15
Figure 3: Learning process of Rainbow-FG with and without intrinsic reward. The two rows have the same meaning with Figure 2.
5 Conclusion
dresses the challenge of sample inefficiency for RL methods in lost-sales IC problems. We design a novel decision framework integrating RLFG and IME to boost the sample efficiency of RL methods. We first specially tailored FG only based on the general property of lost-sales IC problems rather than strong heuristics to generate side experiences to aid RL update. We then conduct a theoretical analysis to demonstrate our method’s effectiveness with Q-learning as an example. The analysis shows that FG decreases the sample complexity by improving the update probabilities for all state-action pairs. Additionally, we design an intrinsic reward to fully utilize FG for lost-sales IC problems based on the analysis result. Experimental results show that our approach greatly enhances RL’s sampling efficiency in IC problems, which is consistent with the theoretical analysis. As for the limitation, we only analyze the magnitude relationship between $\mu_{\text{min}}$ and $\tilde{\mu}_{\text{min}}$; a detailed multiplicative relationship still needs to be proved.